# Natal kick by early-asymmetrical pairs of jets to the neutron star of supernova remnant S147

### v1

@@ 1.1 Two alternative explosion mechanisms
In recent years, studies have heavily referred to two theoretical alternative core-collapse supernova (CCSN) explosion mechanisms: the delayed neutrino explosion mechanism and the jittering-jets explosion mechanism (JJEM). In the delayed neutrino explosion mechanism of CCSNe, neutrinos that the cooling newly born neutron star (NS) emit, heat the gas in the gain region, leading to the explosion (e.g., Andresen et al. 2024; Boccioli & Fragione 2024; Burrows et al. 2024; Janka & Kresse 2024; van Baal et al. 2024; Wang & Burrows 2024; Bamba et al. 2025; Boccioli et al. 2025; Eggenberger Andersen et al. 2025; Huang et al. 2025; Imasheva et al. 2025; Laplace et al. 2025; Maltsev et al. 2025; Maunder et al. 2024; Müller et al. 2025; Nakamura et al. 2025; Sykes & Müller 2025; Janka 2025; Paradiso & Coughlin 2025; Wang & Burrows 2025, for some very recent studies). The gain region is immediately inner to the stalled shock at . Strong turbulence and instabilities in the gain region are crucial to a successful explosion in the neutrino-driven mechanism (e.g., Buellet et al. 2023; Boccioli et al. 2025). The magnetorotational explosion mechanism involves the launching of jets along a fixed axis. Therefore, it requires a rapidly rotating pre-collapse core to launch a pair of jets along a fixed axis (e.g., Shibagaki et al. 2024; Zha et al. 2024; Shibata et al. 2025, for recent studies of this mechanism). The demand for rapid pre-collapse rotation implies that this mechanism is rare; therefore, this mechanism attributes most CCSNe to the neutrino-driven mechanism. We, therefore, consider the magnetorotational explosion mechanism part of the neutrino-driven mechanism.
@@
According to the JJEM, jets explode all CCSNe. In most cases, these are jittering jets. In rare cases, where the core rapidly rotates, the jets are along a fixed axis; some fixed-axis explosions end with a black hole remnant. The exploding, jittering jets come in pairs of opposite jets with fully or partially stochastic direction variations. Intermittent accretion disks or belts around the newly born NS, or a black hole if the NS collapses into a black hole, launch pairs of jets on a time scale of (e.g., Soker 2024a); later jets are possible. The jet-launching process might start as early as after shock bounce, and the two jets might substantially differ in their power and opening angle (Bear et al., 2025); shock bounce is when the shock formation at the center occurs. The source of the stochastic angular momentum of accreted material is instabilities above the NS, seeded by angular momentum fluctuations due to the convection in collapsing core material (e.g., Gilkis & Soker 2014, 2016; Shishkin & Soker 2021, 2023; Wang et al. 2024). These instabilities below the stalled shock at include strong convection and turbulence (e.g., Abdikamalov et al. 2016; Kazeroni & Abdikamalov 2020) and the spiral standing accretion shock instability (e.g., Buellet et al. 2023 on the spiral mode of this instability). Although neutrino heating plays a role in adding energy to the jets (Soker, 2022a), jittering jets are the primary driver of the explosion process and supply most of the explosion energy, even when an energetic magnetar supplies extra post-explosion energy (e.g., Soker & Gilkis 2017; Kumar 2025).
@@
Soker (2024b) summarizes the status of the debate between the two explosion mechanisms at the end of 2024. The neutrino-driven mechanism encounters some small to moderate problems and one major difficulty: (1) Not supplying the explosion energy of energetic CCSNe with explosion energies of , as some CCSNE are observed to have (e.g., Moriya et al. 2025); (2) Predicting somewhat too massive NSs (Soker, 2024b); (3) Predicting failed CCSNe. Recent studies suggest that there is only a small population of failed CCSNe, or not at all (e.g., Byrne & Fraser 2022; Strotjohann et al. 2024; Beasor et al. 2025; Healy et al. 2024). It might be that the neutrino-driven mechanism can overcome this difficulty, as Boccioli et al. (2025) suggested recently. (4) A major problem with the neutrino-driven mechanism is explaining the point-symmetric morphologies of CCSN remnants (CCSNRs); this problem likely rules out the neutrino-driven mechanism as the primary explosion mechanism.111Another noticeable problem of the neutrino-driven mechanism, though, is the qualitative and quantitative disagreements between simulations of different research groups (e.g., Janka (2025)). On the other hand, the JJEM predicts that many CCSNRs exhibit point-symmetric morphologies, and new hydrodynamical simulations demonstrate this (Braudo et al., 2025). The JJEM does not encounter these problems.
@@ 1.2 Bipolar and point-symmetry in CCSNRs
According to the JJEM, launching pairs of opposite jets might shape pairs of structural features in the CCSNR. In general, the jets are not relativistic, as most CCSNe do not have relativistic jets, e.g., Guetta et al. (2020); in the JJEM, relativistic jets, as in gamma ray bursts (e.g., Izzo et al. 2019; Abdikamalov & Beniamini 2025), might be a part of more pairs of jets that explode the star and power the supernova. Since the symmetry axis of the jets in the JJEM changes from one pair to the next, the pairs of structural features in the CCSNR will not share the same axis. This is a point symmetric morphology. The opposite structural features include clumps, filaments, bubbles (faint structures closed and encircled by a brighter rim), lobes (bubbles with partial rims), and ears (protrusions from the main CCSNR shell with decreasing cross sections away from the center). Not all pairs of jets leave morphological imprints, as some are choked deep in the core and are too weak to leave observable marks. Only a fraction of the pairs of exploding jets will leave observable marks. Instabilities in the explosion process, interaction with a circumstellar material (CSM) that the progenitor of the CCSN has lost before explosion (e.g., Chiotellis et al. 2021, 2024; Velázquez et al. 2023; Meyer et al. 2022, 2024b), and interaction with the interstellar medium (e.g., Wu & Zhang 2019; Yan et al. 2020; Lu et al. 2021; Meyer et al. 2024a) cannot form point-symmetric morphologies as Soker & Shishkin (2025) argued. On the contrary, these processes, as well as an NS natal kick, a pulsar wind nebula (PWN), and heating processes such as reverse shock and radioactive decay, are likely to smear the point-symmetrical morphology.
The following is a list of CCSNRs and the studies that attributed their point-symmetric morphologies to the JJEM: SNR 0540-69.3 (Soker, 2022b), CTB 1 (Bear & Soker, 2023), the Vela CCSNR (Soker 2023a; Soker & Shishkin 2025), the Cygnus Loop (Shishkin et al., 2024), N63A (Soker, 2024c), SN 1987A (Soker, 2024d, e), G321.3–3.9 (Soker, 2024f; Shishkin & Soker, 2025a), G107.7-5.1 (Soker, 2024f), W44 (Soker, 2024g), Cassiopeia A (Bear & Soker, 2025), Puppis A (Bear et al., 2025), SNR G0.9+0.1 (Soker, 2025), and the Crab Nebula (Shishkin & Soker, 2024, 2025b).
@@ 1.3 The kick-BEAP mechanism for NS kick velocity
Observations show anti-correlation between the primary X-ray emitting ejecta and the neutron star velocity in Cassiopeia A (e.g., Hwang & Laming 2012). This is compatible with both the tag-boat mechanism and the kick-BEAP mechanism. The asymmetrical ejecta of Puppis A was our motivation in Bear et al. (2025) to propose the kick-BEAP mechanism. In the JJEM, both mechanisms can operate: the kick-BEAP earlier in the explosion, while the tug-boat mechanism operates at relatively later times, (e.g., Wongwathanarat et al. 2013).
@@
The kick-BEAP mechanism can account for the new findings of Valli et al. (2025) who thoroughly analyzed 23 Be X-ray binaries, i.e., with an NS accreting mass from a Be star. Fitting progenitors to the binary eccentricity and orbital period, they identified two populations of NS according to their natal kick velocity: one of NSs with natal kick velocities and one with natal kick velocities around and aligned within from the binary angular momentum. The explosions that formed these NSs were stripped-envelope CCSNe, implying that the progenitors lost large amounts of mass. The mass loss was due to the binary interaction that likely spun up the progenitor. According to the JJEM, the explosion launched one or more powerful pairs of unequal jets along the angular momentum axis. This unequal pair imparted the natal kick to the NS. We find the new findings of Valli et al. (2025) to support the kick-BEAP mechanism in the frame of the JJEM.
@@ 2.1 The general morphology of SNR S147
The CCSNR S147 (Gaze & Shajn, 1952) presents an inhomogeneous structure in the X-ray (e.g., Michailidis et al. 2024; Khabibullin et al. 2024), radio (e.g., Xiao et al. 2008, Khabibullin et al. 2024), and visible (e.g., Drew et al. (2005); Ren et al. (2018)). In emission, e.g., Lozinskaia (1976), it presents many filaments (hence termed the “Spaghetti Nebula”). Dinçel et al. (2015) discovered an OB runaway star inside S147, and suggested the progenitor was in a binary system. CCSNR S147 poses some puzzles, like the kinematics of the NS (PSR J0538+2817) that suggests a younger system age than the estimated age of the SNR, e.g., Khabibullin et al. (2024), who also noted that an expansion of the SNR inside a low-density wind-blown bubble (e.g., Hughes 1987) might solve the age discrepancy. We focus on the relationship between several directions defined by SNR S147: the main jet axis, which we find to likely be two pairs of jets, the line connecting the two X-ray bright regions, and the NS’s proper motion direction. In the Appendix, we use the morphology we identify to estimate the age of S147.
The [O iii] and Hα images of S147 reveal two opposite wide ears, with a filamentary structure that extends into the ears. We present a composite image of these two emission lines in the upper panel of Figure 1, adapted from Mr. Christian Koll. Some earlier studies have connected these two ears with a single line, roughly through the center of the SNR (e.g., Gvaramadze 2006; Grichener & Soker 2017). In the lower panels of Fig 1 we present a close-up of the filamentary structure of the two ears at the edges of S147, from a mosaic within the combined IGAPS survey (Greimel et al., 2021), available on the data website222http://www.star.ucl.ac.uk/IGAPS/data.shtml as imaged in the IPHAS DR2 (Barentsen et al., 2014). We overexpose the images to highlight the sub-structures that define the protrusions and their separation into two pairs. We mark the same dashed lines as in the upper panel for reference.
@@
By examining the possible intersection points between the lines connecting opposite ears from each pair, we can estimate the distribution of intersection points along the NS trajectory. We derive the trajectory using the proper motion estimates of Chatterjee et al. (2009): (see also Ng et al. 2007). The intersection point is the origin of the NS and jets (and hence, explosion), and corresponds to an age of , where the bounds originate from the distribution of points around the mean intersection point. We elaborate on this analysis, our age estimate and previous age estimates in Appendix A.
@@ 2.2 The NS kick episodes
In Figure 2c we show a processed version of the H mosaic from the IGAPS combined survey (Greimel et al., 2021). We pass the original mosaic through a filter (‘Top-hat’ morphological transformation with a 20x20 pixel kernel, as implemented in Bradski 2000) to strengthen sharp edges, as to enhance the filamentary structure defining the general shape and edges of S147. We mark with a yellow-solid line the approximate symmetry axis of the two ears, as we did in Figure 1.
In Figure 2d we combine the X-ray (RGB, Michailidis et al. 2024) and (Greimel et al., 2021) images. We mark the axes from panels (b) and (c). With a dashed cyan line we draw a shifted X-ray axis in the direction of the NS to account for the missing counts from region J (panels (a), (b)), so it starts at the far end of the bright X-ray emission at the north-east (panels (a), (d)). From the current NS location, a red arrow denotes the proper motion direction of Chatterjee et al. (2009). We extend the proper motion direction to form a trajectory and an origin, assuming an age of (Yao et al., 2021), where we mark the trajectory with a dashed red line and the origin with a red dot. For reference, a blue dot marks the geometric center as derived in Kramer et al. (2003). With a broken dotted cyan line, we connect the edges of the shifted X-ray axis to the intersection point of the ear-axis and the NS trajectory, and produce a “bent axis” that signifies an asymmetric pair of jets. Namely, the two opposite jets were at an angle of rather than . We take the directions defined by the two axes (X-ray - cyan, ear-pair - yellow) to be the direction of two NS kick episodes, as marked by the corresponding dashed arrows, and build a vector sum to give the final NS kick velocity (the red-dashed arrow). The ratio of the lengths of the two arrows is , and the angle between them is .
@@ 3 Implications to gravitational waves
The jet pairs that impart the NS kick velocity of SNR S147, and possibly some other CCSNe, are very energetic, and can contribute a fraction of of the explosion energy. We therefore scale up by a factor of five the energy of the predicted gravitational wave signal from the most energetic jet-inflated bubbles in the JJEM. We present the results in Figure 3. We adapted a figure from Mezzacappa et al. (2023), and added the predicted characteristic of the gravitational wave with a horseshoe-shaped yellow zone, scaled from Soker (2023b). The noisy blue line represents a gravitational wave prediction based on the neutrino-driven mechanism. As the turbulent motion near the newly born neutron star is similar in both explosion mechanisms, we predict the total gravitational wave signal to be the combination of the noisy blue line and the horseshoe-shaped yellow zone.
Gravitational waves can teach us about some CCSN explosion properties (e.g., Mitra et al. 2024); hence, we should include all possible CCSN sources of gravitational waves. Soker (2023b) concluded that detecting gravitational waves from jet-inflated bubbles is possible for a Galactic CCSN, but marginal. We conclude that a clear signal will be detected from a galactic CCSN, possibly even a CCSN in the Magellanic Clouds.

### v2

@@ Abstract
show]s.dmitry@campus.technion.ac.il
]ealeal44@technion.ac.il
]soker@physics.technion.ac.il
@@ 1.1 Two alternative explosion mechanisms
In recent years, studies have heavily referred to two theoretical alternative core-collapse supernova (CCSN) explosion mechanisms: the delayed neutrino explosion mechanism and the jittering-jets explosion mechanism (JJEM). In the delayed neutrino explosion mechanism of CCSNe, neutrinos that the cooling newly born neutron star (NS) emit, heat the gas in the ‘gain region’, leading to the explosion (e.g., Andresen et al. 2024; Boccioli & Fragione 2024; Burrows et al. 2024; Janka & Kresse 2024; van Baal et al. 2024; Wang & Burrows 2024; Bamba et al. 2025; Boccioli et al. 2025; Eggenberger Andersen et al. 2025; Huang et al. 2025; Imasheva et al. 2025; Janka 2025; Laplace et al. 2025; Maltsev et al. 2025; Maunder et al. 2024; Mori et al. 2025; Müller et al. 2025a; Nakamura et al. 2025; Sykes & Müller 2025; Paradiso & Coughlin 2025; Vink et al. 2025; Wang & Burrows 2025, for some very recent studies supporting the neutrino-driven mechanism). The gain region is immediately inner to the stalled shock at . Strong turbulence and instabilities in the gain region are crucial to a successful explosion in the neutrino-driven mechanism (e.g., Buellet et al. 2023; Boccioli et al. 2025). The magnetorotational explosion mechanism involves the launching of jets along a fixed axis. Therefore, it requires a rapidly rotating pre-collapse core to launch a pair of jets along a fixed axis (e.g., Shibagaki et al. 2024; Zha et al. 2024; Shibata et al. 2025, for recent studies of this mechanism). The demand for rapid pre-collapse rotation implies that this mechanism is rare; therefore, this mechanism attributes most CCSNe to the neutrino-driven mechanism. We, therefore, consider the magnetorotational explosion mechanism part of the neutrino-driven mechanism.
@@
According to the JJEM, jets explode all CCSNe. In most cases, these are jittering jets. In rare cases, where the core rapidly rotates, the jets are along a fixed axis; some fixed-axis explosions end with a black hole remnant. The exploding, jittering jets come in pairs of opposite jets with fully or partially stochastic direction variations. Intermittent accretion disks or belts around the newly born NS, or a black hole if the NS collapses into a black hole, launch pairs of jets on a time scale of (e.g., Soker 2025a); later jets are possible. The jet-launching process might start as early as after shock bounce, and the two jets might substantially differ in their power and opening angle (Bear et al., 2025); shock bounce is when the shock formation at the center occurs. The source of the stochastic angular momentum of accreted material is instabilities above the NS, seeded by angular momentum fluctuations due to the convection in collapsing core material (e.g., Gilkis & Soker 2014, 2016; Shishkin & Soker 2021, 2023; Wang et al. 2024). These instabilities below the stalled shock at include strong convection and turbulence (e.g., Abdikamalov et al. 2016; Kazeroni & Abdikamalov 2020) and the spiral standing accretion shock instability (e.g., Buellet et al. 2023 on the spiral mode of this instability). Although neutrino heating plays a role in adding energy to the jets (Soker, 2022a), jittering jets are the primary driver of the explosion process and supply most of the explosion energy, even when an energetic magnetar supplies extra post-explosion energy (e.g., Soker & Gilkis 2017; Kumar 2025).
@@
Soker (2024a) summarized the status of the debate between the two explosion mechanisms at the end of 2024, The neutrino-driven mechanism encounters some small to moderate problems and one major difficulty: (1) Not supplying the explosion energy of energetic CCSNe with explosion energies of , as some CCSNE are observed to have (e.g., Moriya et al. 2025); (2) Predicting NS mass distribution that is displaced to somewhat higher NS masses than observed (Soker 2024a). (3) Predicting failed CCSNe. Recent studies suggest that there is only a small population of failed CCSNe, or not at all (e.g., Byrne & Fraser 2022; Strotjohann et al. 2024; Beasor et al. 2025; Healy et al. 2024). It might be that the neutrino-driven mechanism can overcome this difficulty, as Boccioli et al. (2025) suggested recently. (4) A major problem with the neutrino-driven mechanism is explaining the point-symmetric morphologies of CCSN remnants (CCSNRs); this problem likely rules out the neutrino-driven mechanism as the primary explosion mechanism.111Another noticeable problem of the neutrino-driven mechanism, though, is the qualitative and quantitative disagreements between simulations of different research groups (e.g., Janka (2025)). On the other hand, the JJEM predicts that many CCSNRs exhibit point-symmetric morphologies, and new hydrodynamical simulations demonstrate this (Braudo et al., 2025). The JJEM does not encounter these problems.
@@ 1.2 Bipolar and point-symmetry in CCSNRs
According to the JJEM, launching pairs of opposite jets might shape pairs of structural features in the CCSNR. In general, the jets are not relativistic, as most CCSNe do not have relativistic jets, e.g., Guetta et al. (2020); in the JJEM, relativistic jets, as in gamma ray bursts (e.g., Izzo et al. 2019; Abdikamalov & Beniamini 2025), might be a part of more pairs of jets that explode the star and power the supernova. Since the symmetry axis of the jets in the JJEM changes from one pair to the next, the pairs of structural features in the CCSNR will not share the same axis. This is a point symmetric morphology. Not all pairs of jets leave morphological imprints, as some are choked deep in the core and are too weak to leave observable marks. Only a fraction of the pairs of exploding jets will leave observable marks. Instabilities in the explosion process, interaction with a circumstellar material (CSM) that the progenitor of the CCSN has lost before explosion (e.g., Chiotellis et al. 2021, 2024; Velázquez et al. 2023; Meyer et al. 2022, 2024b), and interaction with the interstellar medium (e.g., Wu & Zhang 2019; Yan et al. 2020; Lu et al. 2021; Meyer et al. 2024a) cannot form point-symmetric morphologies as Soker & Shishkin (2025) argued. On the contrary, these processes, as well as an NS natal kick, a pulsar wind nebula (PWN), and heating processes such as reverse shock and radioactive decay, are likely to smear the point-symmetrical morphology.
The following is a list of CCSNRs and the studies that attributed their point-symmetric morphologies to the JJEM: SNR 0540-69.3 (Soker, 2022b), CTB 1 (Bear & Soker, 2023), the Vela CCSNR (Soker 2023a; Soker & Shishkin 2025), the Cygnus Loop (Shishkin et al., 2024), N63A (Soker, 2024b), SN 1987A (Soker, 2024c, d), G321.3–3.9 (Soker, 2024e; Shishkin & Soker, 2025a), G107.7-5.1 (Soker, 2024e), W44 (Soker, 2024f), Cassiopeia A (Bear & Soker, 2025), Puppis A (Bear et al., 2025), SNR G0.9+0.1 (Soker, 2025b), N132D (Soker, 2025c), Circinus X-1 (Soker & Akashi, 2025), and the Crab Nebula (Shishkin & Soker, 2024, 2025b).
@@ 1.3 The kick-BEAP mechanism for NS kick velocity
Observations show anti-correlation between the primary X-ray emitting ejecta and the neutron star velocity in Cassiopeia A (e.g., Hwang & Laming 2012). This is compatible with both the tug-boat mechanism and the kick-BEAP mechanism. The asymmetrical ejecta of Puppis A was our motivation in Bear et al. (2025) to propose the kick-BEAP mechanism. In the JJEM, both mechanisms can operate: the kick-BEAP earlier in the explosion, while the tug-boat mechanism operates at relatively later times, (e.g., Wongwathanarat et al. 2013).
@@
The kick-BEAP mechanism can account for the new findings of Valli et al. (2025) who thoroughly analyzed 23 Be X-ray binaries, i.e., with an NS accreting mass from a Be star. Fitting progenitors to the binary eccentricity and orbital period, they identified two populations of NS according to their natal kick velocity: one of NSs with natal kick velocities and one with natal kick velocities around and aligned within from the binary angular momentum. The explosions that formed these NSs were stripped-envelope CCSNe, implying that the progenitors lost large amounts of mass. The mass loss was due to the binary interaction that likely spun up the progenitor (e.g., Gilkis et al. 2025). According to the JJEM, the explosion launched one or more powerful pairs of unequal jets along the angular momentum axis. This unequal pair imparted the natal kick to the NS. We find the new findings of Valli et al. (2025) to support the kick-BEAP mechanism in the frame of the JJEM.
@@ 2.1 The general morphology of SNR S147
The CCSNR S147 (Gaze & Shajn, 1952) presents an inhomogeneous structure in the X-ray (e.g., Michailidis et al. 2024; Khabibullin et al. 2024b), radio (e.g., Xiao et al. 2008, Khabibullin et al. 2024a), and visible (e.g., Drew et al. 2005; Ren et al. 2018). In H emission, e.g., Lozinskaia (1976), it presents many filaments (hence termed the “Spaghetti Nebula”). Dinçel et al. (2015) discovered an OB runaway star inside S147, and suggested the progenitor was in a binary system. CCSNR S147 poses some puzzles, like the kinematics of the NS (PSR J0538+2817) that suggests a younger system age than the estimated age of the SNR, e.g., Khabibullin et al. (2024b), who also noted that an expansion of the SNR inside a low-density wind-blown bubble (e.g., Hughes 1987) might solve the age discrepancy. We focus on the relationship between several directions defined by SNR S147: the main jet axis, which we find to likely be two pairs of jets, the line connecting the two X-ray bright regions, and the NS’s proper motion direction. In the Appendix, we use the morphology we identify to estimate the age of S147.
The [O iii] and H images of S147 reveal two opposite wide ears, with a filamentary structure that extends into the ears. We present a composite image of these two emission lines in the upper panel of Figure 1, adapted from Mr. Christian Koll. Some earlier studies have connected these two ears with a single line, roughly through the center of the SNR (e.g., Gvaramadze 2006; Grichener & Soker 2017). In the lower panels of Fig 1 we present a close-up of the H filamentary structure of the two ears at the edges of S147, from a mosaic within the combined IGAPS survey (Greimel et al., 2021), available on the data website222http://www.star.ucl.ac.uk/IGAPS/data.shtml as imaged in the IPHAS DR2 (Barentsen et al., 2014). We overexpose the images to highlight the sub-structures that define the protrusions and their separation into two pairs. We mark the same dashed lines as in the upper panel for reference.
@@
By examining the possible intersection points between the lines connecting opposite ears from each pair, we can estimate the distribution of intersection points along the NS trajectory. We derive the trajectory using the proper motion estimates of Chatterjee et al. (2009): (see also Ng et al. 2007). The intersection point is the origin of the NS and jets (and hence, explosion), and corresponds to an age of , where the bounds originate from the distribution of points around the mean intersection point. We elaborate on this analysis, , our age estimate and previous age estimates in Appendix A.
@@ 2.2 The NS kick episodes
In Figure 2c we show a processed version of the H mosaic from the IGAPS combined survey (Greimel et al., 2021). We pass the original mosaic through a filter (‘Top-hat’ morphological transformation with a 20x20 pixel kernel, as implemented in ; Bradski 2000) to strengthen sharp edges, as to enhance the filamentary structure defining the general shape and edges of S147. We mark with a yellow-solid line the approximate symmetry axis of the two ears, as we did in Figure 1.
In Figure 2d we combine the X-ray (RGB, Michailidis et al. 2024) and H (Greimel et al., 2021) images. We mark the axes from panels (b) and (c). With a dashed cyan line we draw a shifted X-ray axis in the direction of the NS to account for the missing counts from region J (panels (a), (b)), so it starts at the far end of the bright X-ray emission at the north-east (panels (a), (d)). We extend the proper motion (Chatterjee et al., 2009) direction to form a trajectory and an origin, assuming an age of (Yao et al., 2021), where we mark the trajectory with a solid red line and the origin with a red dot. For reference, a blue dot marks the geometric center as derived in Kramer et al. (2003). With a broken dotted cyan line, we connect the edges of the shifted X-ray axis to the intersection point of the ear-axis and the NS trajectory, and produce a “bent axis” that signifies an asymmetric pair of jets. Namely, the two opposite jets were at an angle of rather than . We take the directions defined by the two axes (X-ray - cyan, ear-pair - yellow) to be the direction of two NS kick episodes, as marked by the corresponding dashed arrows, and build a vector sum to give the final NS kick velocity (the red-solid arrow). The ratio of the lengths of the two arrows is , and the angle between them is .
@@ 3 Implications to gravitational waves
The jet pairs that impart the NS kick velocity of SNR S147, and possibly some other CCSNe, are very energetic, and can contribute a fraction of of the explosion energy. We, therefore, scale up by a factor of five the energy of the predicted gravitational wave signal from the most energetic jet-inflated bubbles in the JJEM. We present the results in Figure 3. We adapted a figure from Mezzacappa et al. (2023), and added the predicted characteristic of the gravitational wave with a horseshoe-shaped yellow zone, scaled from Soker (2023b). The noisy blue line represents a gravitational wave prediction based on the neutrino-driven mechanism. As the turbulent motion near the newly born neutron star is similar in both explosion mechanisms, we predict the total gravitational wave signal to be the combination of the noisy blue line and the horseshoe-shaped yellow zone.
Gravitational waves can teach us about some CCSN explosion properties (e.g., Iess et al. 2023; Mitra et al. 2024); hence, we should include all possible CCSN sources of gravitational waves. Soker (2023b) concluded that detecting gravitational waves from jet-inflated bubbles is possible for a Galactic CCSN, but marginal. We conclude that a clear signal will be detected from a galactic CCSN, possibly even a CCSN in the Magellanic Clouds.